OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). If necessary, use the slash mark (/) for the division symbol.

OpenStudy (anonymous):

\[F(x)=\frac{ 2x }{ 7+4 }\]

OpenStudy (anonymous):

\[f ^{-1}(y)=\]

OpenStudy (anonymous):

@Luigi0210 @Anickyan i really lost on this problem

OpenStudy (luigi0210):

Oh, fun one :)

OpenStudy (anonymous):

\(y = \frac{2x}{7 + 4}\) solve for \(x\)

OpenStudy (anonymous):

x+2/11?

OpenStudy (anonymous):

Nope. \(2x\) means \(2 * x\)

OpenStudy (anonymous):

so 2+x/11

OpenStudy (anonymous):

is the^ right

OpenStudy (anonymous):

No. Do you want me to keep on explaining or give you the answer?

OpenStudy (anonymous):

@Anickyan

OpenStudy (anonymous):

what very u think is best for me

OpenStudy (anonymous):

I think explaining it the best thing. I can give you some easy tasks to do: http://www.sosmath.com/algebra/solve/solve0/solve0.html They start easy, and get more and more difficult.

OpenStudy (anonymous):

damn thats really hard

OpenStudy (anonymous):

hey can u help

OpenStudy (anonymous):

@Anickyan @Luigi0210

OpenStudy (luigi0210):

@Anickyan Already explained it..

OpenStudy (anonymous):

i keep on getting 2+x/11

OpenStudy (anonymous):

i give up its hard

OpenStudy (anonymous):

@mathstudent55 can u check mt answer

OpenStudy (debbieg):

You have \(F(x)=\dfrac{ 2x }{ 7+4 }\) which is really just \(F(x)=\dfrac{ 2x }{ 11 }\) Replace the F(x) with y: \(y=\dfrac{ 2x }{ 11 }\) Now solve it for x.

OpenStudy (debbieg):

You should NOT get getting \(y=2+\dfrac{ x }{ 11 }\) In fact, you should be getting x={some stuff involving y|

OpenStudy (kropot72):

The inverse of f(x) should be written as \[f ^{-1} (x)\]

OpenStudy (kropot72):

So the question is stated incorrectly.

OpenStudy (debbieg):

@kropot72 , I agree but we had a big thread about this yesterday, lol. @romanortiz65 's instructor apparently uses different notation.

OpenStudy (anonymous):

x-2/11?

OpenStudy (debbieg):

@kropot72 this will explain it: http://openstudy.com/study#/updates/521b8fdbe4b06211a67d369e

OpenStudy (debbieg):

@romanortiz65 you have \(y=\dfrac{ 2x }{ 11 }\) And you need to solve for x. what is STEP 1. Tell the first thing you will do.

OpenStudy (anonymous):

switch y and x

OpenStudy (debbieg):

That isn't how you "solve for x". You solve for x by isolating x, getting it all alone.

OpenStudy (anonymous):

oh x+2/11

OpenStudy (debbieg):

Again, you should have x ALONE. and it should be = something. I'm not sure why you are giving responses that are not equations, and that don't isolate x. x is not involved in any sums/differences, so it isn't likely that you need any sums/differences in the inverse.

OpenStudy (debbieg):

Look at the equation this way: \(y=\dfrac{ 2x }{ 11 }\) is the same as \(y=x\cdot\dfrac{ 2 }{ 11 }\) Now how do you solve that for x? How can you get that x= all alone on the RHS of the equation?

OpenStudy (debbieg):

HINT: IF \(t=r\cdot\dfrac{ 5 }{ 13 }\) THEN \(\dfrac{ 13 }{ 5 }\cdot t=r\)

OpenStudy (anonymous):

-x -x both sides

OpenStudy (anonymous):

@DebbieG

OpenStudy (anonymous):

y=13/5?

OpenStudy (anonymous):

@Luigi0210 check answer

OpenStudy (anonymous):

@hba check anwer please

OpenStudy (luigi0210):

Where are your variables..?

OpenStudy (anonymous):

can u just please just give me the answer im running super late for school please

OpenStudy (debbieg):

You have \(y=x\cdot\dfrac{ 2 }{ 11 }\) Read my hint above. You can multiply both sides by the reciprocal of 2/11:

OpenStudy (debbieg):

\(y=x\cdot\dfrac{ 2 }{ 11 }\) \(\dfrac{ 11 }{ 2 }\cdot y=x\cdot\dfrac{ 2 }{ 11 }\cdot\dfrac{ 11 }{ 2 }\) What do you get?

OpenStudy (debbieg):

@incognito, if you read above and follow the link to the prior question, you'll see the explanation about his instructor wanting the inverse as a function of y, not of x.

OpenStudy (anonymous):

x x 22/22

OpenStudy (debbieg):

@romanortiz65 I don't even know what you mean by that. where is your y?? Where is your = sign?

OpenStudy (anonymous):

just give me final answer running late for my senior picture

OpenStudy (anonymous):

please

OpenStudy (debbieg):

Oh, ok.... I think you mean that's the RHS?? I guess so. But not sure why you have an extra x.

OpenStudy (luigi0210):

Dude.. it's just one question. It's not like you're gonna die or fail if you don't answer it.

OpenStudy (debbieg):

I pretty much did. Just simplify that RHS above and you will have x=???

OpenStudy (anonymous):

y=22/22

OpenStudy (luigi0210):

Well then..

OpenStudy (debbieg):

HOW does.... \(\dfrac{ 11 }{ 2 }\cdot y=x\cdot\dfrac{ 2 }{ 11 }\cdot\dfrac{ 11 }{ 2 }\) Simplify to y=22/22??? 1. Where did the fraction go from the LHS?? 2. Where did the x go from the RHS?? 3. 22/22 is not simplified. It can be made simpler!! C'mon, man, you are psyching yourself out. You're so convinced that you can't do this that you aren't thinking it through.

OpenStudy (anonymous):

11/11

OpenStudy (anonymous):

1

OpenStudy (debbieg):

\[\dfrac{\cancel 2 }{ \cancel {11} }\cdot\dfrac{\cancel {11} }{\cancel 2 }=1\]

OpenStudy (anonymous):

i got this one wrong ;( didnt pass my test

OpenStudy (anonymous):

@DebbieG the ?s u helped me were wrongg